OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (–2, –3) and has a slope of

OpenStudy (hoblos):

of ...

OpenStudy (eyust707):

so standard form is y= mx +b

OpenStudy (eyust707):

m will be your slope.. if we have a slope and two points we can use them to find b

OpenStudy (anonymous):

1/2

OpenStudy (eyust707):

okay so now we have y=(1/2)x +b

OpenStudy (eyust707):

whats is one of your y's equal to?

OpenStudy (eyust707):

and what about one of your x's?

OpenStudy (anonymous):

I thought standard form is a(x)+b(x)=c

OpenStudy (eyust707):

ahh yes you are right

OpenStudy (anonymous):

yeah I don't know how to do this at all

OpenStudy (eyust707):

okay we can still use what we have done

OpenStudy (eyust707):

so y=(1/2)x +b and since we know that at one point y=-3 and x=-2 we can plug those in to find b

OpenStudy (eyust707):

-3= (1/2)(-2) +b -3=-1 +b -2 = b

OpenStudy (eyust707):

thats our b for all points

OpenStudy (eyust707):

y=(1/2)x -2

OpenStudy (anonymous):

that's what I ended up with but what would it be in standard form?

OpenStudy (eyust707):

now just move the x term over

OpenStudy (eyust707):

y - (1/2) x =-2

OpenStudy (anonymous):

here's the answers I was given.. 2x + y = 4 2x – y = 4 x + 2y = 4 x – 2y = 4

OpenStudy (eyust707):

we can simplfy our answer to look like theres if you would like, just multiply every term by -2

OpenStudy (anonymous):

damn really..I didn't even think of that thank you so much!

OpenStudy (eyust707):

yes i guess they are a little picky with that x term being positive... so you should always do that at the end

OpenStudy (eyust707):

no problem... math is easy once ya get it!